### 17 Dec 2023 v1

# Policy Optimization in RLHF: The Impact of Out-of-preference Data

Abstract
Aligning intelligent agents with human preferences and values is important. This paper examines two popular alignment methods: Direct Preference Optimization (DPO) and Reward-Model-Based Policy Optimization (RMB-PO). A variant of RMB-PO, referred to as RMB-PO+ is also considered. These methods, either explicitly or implicitly, learn a reward model from preference data and differ in the data used for policy optimization to unlock the generalization ability of the reward model. In particular, compared with DPO, RMB-PO additionally uses policy-generated data, and RMB-PO+ further leverages new, preference-free data. We examine the impact of such out-of-preference data. Our study, conducted through controlled and synthetic experiments, demonstrates that DPO performs poorly, whereas RMB-PO+ performs the best. In particular, even when providing the policy model with a good feature representation, we find that policy optimization with adequate out-of-preference data significantly improves performance by harnessing the reward model’s generalization capabilities. [^1] Code is available at https://github.com/liziniu/policy_optimization
1 Introduction
Developing intelligent agents requires alignment with human values (Russell and Norvig, 2010). A standard practice involves providing a human preference dataset for the agent to learn from. According to utility theory (Fishburn et al., 1979), preference is connected with a certain reward function. Currently, there are two kinds of alignment methods: the reward-model-free approach, such as Direct Preference Optimization (DPO) (Rafailov et al., 2023) and Identity Policy Optimization (IPO) (Azar et al., 2023), and the reward-model-based approaches, exemplified by the so-called Reinforcement Learning from Human Feedback (RLHF) (Christiano et al., 2017; Stiennon et al., 2020; OpenAI, 2023). Specifically, the first method directly learns a decision policy from a preference dataset, without the need to train a separate reward model. On the other hand, the second approach involves training a reward model from the preference data and subsequently employing policy optimization (e.g., reinforcement learning) on new, preference-free data. While both approaches have been shown to improve the decision policy by leveraging preference data, the superiority of one method over the other remains unclear, a determination crucial for future advancements.
We briefly explain why this question is hard to answer. The main difficulty lies in the intuition that the policy model, powered by a (pre-trained) neural network or learned from preference-data, possesses a certain generalization ability. Therefore, it is unclear whether policy optimization on additional data, with an imperfect yet generalizable reward model, is necessary.
In this paper, we explore the above question through controlled synthetic experiments. We study two contextual bandit tasks with linear function approximation and neural function approximation, respectively. Our studies take a stochastic optimization viewpoint. Concretely, we argue that the policy optimization in these methods corresponds to different versions of stochastic approximation of the (expected) reward maximization problem. In particular, with a reward model, stochastic approximation can be performed more accurately with out-of-preference data, which interestingly does not increase the sample complexity of preference data.
Our experiments validate the above ideas. One main experiment, where we manually ensure that the policy shares the same good feature representation with the reward model, shows that policy optimization with additional out-of-preference data still improves generalization performance. Other experiments also support this claim. Finally, we provide a discussion about this phenomenon with reference to other fields, such as imitation learning and reinforcement learning.
2 Problem Formulation
We consider the so-called contextual bandit (Langford and Zhang, 2007; Lu et al., 2010) formulation for the alignment problem. Let $s$ and $a$ be the state (i.e., context) and decision action, respectively. We aim to obtain a decision policy $\pi$ that acts optimally in terms of reward maximization:
$\pi^{\star}\leftarrow\mathop{\rm argmax}_{\pi}\mathbb{E}_{s\sim\rho(\cdot)}\mathbb{E}_{a\sim\pi(\cdot|s)}\left[r(s,a)\right],$ (1)
where the symbol $\rho$ denotes the state distribution, and $r$ is the ground truth reward function. In the context of alignment, the difficulty is that the reward function is unknown but only preferences over two actions are observed. Typically, the Bradley-Terry-Luce (Bradley and Terry, 1952) assumption is used:
${\mathbb{P}}\left(a>a^{\prime}|s\right)=\frac{\exp(r(s,a))}{\exp(r(s,a))+\exp(r(s,a^{\prime}))},$
where the symbol $a>a^{\prime}$ means that $a$ is more preferred compared with $a^{\prime}$. Given a preference dataset $D_{\operatorname{pref}}=\{(s_{i},a_{i},a_{i}^{\prime})\}_{i=1}^{n}$, where $a_{i}>a_{i}^{\prime}$ is assumed without loss of generality, the reward learning objective is
$\widehat{r}\leftarrow\mathop{\rm argmax}_{r}\sum_{i=1}^{n}\log\left(\sigma\left(r(s_{i},a_{i})-r(s_{i},a_{i}^{\prime})\right)\right),\quad(s_{i},a_{i},a_{i}^{\prime})\sim D_{\operatorname{pref}}.$ (2)
where $\sigma(\cdot)$ is the sigmoid function. Let $\pi_{\operatorname{ref}}$ be a reference policy model and $\beta>0$ be a hyper-parameter. The Reward-Model-Based Policy Optimization (RMB-PO) approach optimizes the policy on the policy-generated data by sampling states from the preference dataset:
$\widehat{\pi}_{\operatorname{RMB-PO}}\leftarrow\mathop{\rm argmax}_{\pi}\sum_{i=1}^{n}\mathbb{E}_{a\sim\pi(\cdot|s_{i})}\left[\widehat{r}(s_{i},a)\right]+\beta D_{\mathrm{KL}}(\pi(\cdot|s_{i}),\pi_{\operatorname{ref}}(\cdot|s_{i})),\quad s_{i}\sim D_{\operatorname{pref}}.$ (3)
In (Ouyang et al., 2022; Touvron et al., 2023), a variant of RMB-PO, refereed to as RMB-PO+ in this paper, further leverages a new, preference-free dataset $D_{\operatorname{new}}=\{s_{j}\}_{j=1}^{m}$ for policy optimization:
$\widehat{\pi}_{\operatorname{RMB-PO+}}\leftarrow\mathop{\rm argmax}_{\pi}\sum_{j=1}^{m}\mathbb{E}_{a\sim\pi(\cdot|s_{j})}\left[\widehat{r}(s_{j},a)\right]+\beta D_{\mathrm{KL}}(\pi(\cdot|s_{j}),\pi_{\operatorname{ref}}(\cdot|s_{j})),\,\,s_{j}\sim D_{\operatorname{new}}.$ (4)
Note that the dataset $D_{\operatorname{new}}$ is cheap to obtain and usually $m\geq n$ (Ouyang et al., 2022).
In addition to the above methods, the Direct Preference Optimization (DPO) is developed in (Rafailov et al., 2023), which combines Equation 2 and Equation 3 to a single supervised objective:
$\widehat{\pi}_{\operatorname{DPO}}\leftarrow\mathop{\rm argmax}_{\pi}\sum_{i=1}^{n}\log\sigma\left(\beta\log\frac{\pi(a_{i}|s_{i})}{\pi_{\operatorname{ref}}(a_{i}|s_{i})}-\beta\log\frac{\pi(a_{i}^{\prime}|s_{i})}{\pi_{\operatorname{ref}}(a_{i}^{\prime}|s_{i})}\right),\quad(s_{i},a_{i},a_{i}^{\prime})\sim D_{\operatorname{pref}}.$ (5)
Note that DPO uses solely the preference data, without policy-generated data or preference-free data.
2.1 A Stochastic Optimization Perspective
Based on the above formulations, we argue that all the mentioned methods are stochastic approximations to Equation 1, with errors arising from three sources:
  • 1)
    the reward model error $|\widehat{r}(s,a)-r(s,a)|$;
  • 2)
    the estimation error when using finite samples to calculate the expectation $\mathbb{E}_{a\sim\pi(\cdot|s)}[\cdot]$;
  • 3)
    the estimation error when using finite samples to calculate the expectation $\mathbb{E}_{s\sim\rho(\cdot)}[\cdot]$.
The first error primarily results from having only finite preference data, which is expensive to obtain. Compared with DPO, RMB-PO aims to mitigate the second error, while RMB-PO+ further reduces the third error. We note that RMB-PO and RMB-PO+ do not increase the sample complexity of preference data but only incur additional computation steps.
In our study, we assume that the action space $\mathcal{A}$ is finite and the expectation $\mathbb{E}_{a\sim\pi(\cdot|s)}[\cdot]$ is easy to obtain. For applications with combinatorial action spaces (e.g., large language models), computing this expectation is generally intractable, and practical algorithms may employ the policy to generate stochastic trials to approximate the expectation; refer to, e.g., (Li et al., 2023). We note that even in this case, the sample complexity of preference data does not increase.
Another viewpoint is that the RMB-PO and RMB-PO+ methods incorporate additional policy-generated data and preference-free data, essentially employing a form of data augmentation. We summarize the effect of such data augmentation in Table 1.
DPO RMB-PO RMB-PO+
Improvement of Stochastic Approximation in State Space
Improvement of Stochastic Approximation in Action Space
Table 1: Illustration of the data augmentation effect in RMB-PO and RMB-PO+.
3 Experiments
In this section, we conduct numerical experiments to validate the improvement of RMB-PO and RMB-PO+ by better stochastic approximation. All of our experiments are run with 10 different random seeds (2021-2030), and the averaged results are reported. Note that we set $\pi_{\text{REF}}$ to be a policy with a uniform action distribution in all experiments and $\beta=0.01$ for all methods. Besides, we use a policy with a uniform action distribution to collect the preference data.
3.1 Linear Bandit
We study a linear bandit task, where we have $r(s,a)=\phi_{r}(s,a)^{\top}\theta_{r}^{\star}$, with $\phi_{r}(s,a)\in\mathbb{R}^{d}$ denoting the feature representation and $\theta_{r}^{\star}\in\mathbb{R}^{d}$ as the parameter. In this case, the reward learning optimization problem is convex, so we use CVXPY (Diamond and Boyd, 2016) to find the solution $\widehat{r}$. In particular, we use the feature map $\phi_{r}(s,a)$ and the parameter $\theta_{r}^{\star}$ as
$\phi_{r}(s,a)=\left((a+1)\cdot\cos(s\cdot\pi),\frac{1}{a+1}\cdot\sin(s\cdot\pi)\right)^{\top},\quad\theta_{r}^{\star}=(1,2)^{\top},$
where $s\in\mathcal{S}=[0,1]$ and $a\in\mathcal{A}=\{0,1,2,3\}$. A uniform distribution over ${\mathcal{S}}$ is studied. For the policy, we consider the parameterization
$\pi(a|s)=\frac{\exp(\phi_{\pi}(s,a)^{\top}\theta_{\pi})}{\sum_{a^{\prime}}\exp(\phi_{\pi}(s,a^{\prime})^{\top}\theta_{\pi})},$
with $\phi_{\pi}(s,a)$ and $\theta_{\pi}$ both in ${}^{2}$. In this case, the policy optimization problem is a non-convex problem, but the gradient domination condition holds (Agarwal et al., 2021). We use the gradient ascent method with the AdaGrad optimizer (Duchi et al., 2011) (a step size of 0.1 is used).
Figure 1: Optimality gap with $\phi_{\pi}=\phi_{r}$.
Figure 1: Optimality gap with $\phi_{\pi}=\phi_{r}$.
Figure 2: Optimality gap with $\phi_{\pi}\neq\phi_{r}$.
We examine two scenarios. In the first scenario, there is no feature mismatch between the reward and policy models, i.e., $\phi_{\pi}=\phi_{r}$. In the second, we use a different feature map for policy:
$\phi_{\pi}(s,a)=\left((a+1)\cdot\sin(s\cdot\pi),\frac{1}{a+1}\cdot\cos(s\cdot\pi)\right)^{\top}.$
In our experiments, we set the size of preference data to be $n=20$ and the size of preference-free data to be $m=10n$, resulting in training accuracy of the reward model ranging from 60% to 80%. We display the optimality gap $|r(\pi^{\star})-r(\widehat{\pi})|$ (the smaller, the better) in Figure 2 and Figure 2, where $r(\pi)$ is the expected performance, i.e., $r(\pi)=\mathbb{E}_{s\sim\rho(\cdot)}\mathbb{E}_{a\sim\pi(\cdot|s)}[r(s,a)]$ (in our experiments, we use 5000 sampled states to approximate this expectation).
From Figure 2, we see that even though the policy model is provided with a good feature (e.g., in Figure 2), RMB-PO methods can benefit from out-of-preference data. In the case where $\phi_{\pi}\neq\phi_{r}$ in Figure 2, we find that RMB-PO+ is better than RMB-PO by leveraging additional preference-free data. Thus, we believe it is crucial to learn the optimal action (as inferred by the reward model) on out-of-preference data, even when the two models share the same good feature.
Following the same setup, we provide ablation studies regarding the size of preference-free data used in RMB-PO+. See the results in Figure 4 and Figure 4. We find that the previous conclusions still hold true.
Figure 3: Optimality gap with $\phi_{\pi}=\phi_{r}$.
Figure 3: Optimality gap with $\phi_{\pi}=\phi_{r}$.
Figure 4: Optimality gap with $\phi_{\pi}\neq\phi_{r}$.
3.2 Neural Bandit
In this section, we study a neural bandit problem. Specifically, we study the case where $r(s,a)=f_{\theta^{\star}_{r}}(s,a)$, with $f_{\theta_{r}^{\star}}$ being a fixed 1-hidden-layer multi-layer perceptron (MLP) neural network, having a hidden size of 64. For reward learning, we use a 2-hidden-layer MLP with a hidden size of 64, and the policy network is also a 2-hidden-layer MLP with a hidden size of 64. We consider a continuous state space $\mathcal{S}=[-1,1]^{50}$ and a discrete action space $\mathcal{A}=\{0,1,2,\ldots,9\}$. The state distribution $\rho$ is uniform and one-hot feature representation for actions is used.
(a) Reward neural network.
(b) Policy neural network.
Figure 5: Architectures of the reward and policy models.
We note that, unlike in the linear bandit case where we could fix the feature representations of the reward and policy models to be the same, in this case, the feature representations of the reward and policy models are purely learned from the given data. The architectures of the reward and policy models are shown in Figure 5. All neural networks are optimized using the Adam optimizer (Kingma and Ba, 2015) with a step size of $10^{-3}$.
We run experiments with varying sizes of preference-free data $m$ while fixing the preference data size at $n=50$. We report the results in Figure 6. First, we observe that RMB-PO and RMB-PO+ significantly outperform DPO. Furthermore, simply using a preference-free data size that is twice as large already improves performance over RMB-PO, and further scaling does not help too much.
Figure 6: Optimality gap of learned policies in the neural bandit task.
4 Discussion
Our results are related to imitation learning (Osa et al., 2018), where the goal is to learn a policy from expert demonstrations. A popular approach to achieve this goal is through behavioral cloning (BC) (Pomerleau, 1991), which trains a policy model by maximizing the likelihood of expert data. Note that the working mechanism of BC is quite similar to DPO, as in Equation 5, where the likelihood of positively preferred actions is increased and that of negatively preferred actions is decreased:
$\pi_{\operatorname{BC}}\leftarrow\mathop{\rm argmax}_{\pi}\sum_{i=1}^{n}\log\pi(a_{i}|s_{i}),\quad(s_{i},a_{i})\sim D_{\operatorname{E}},$
where $D_{\operatorname{E}}$ is the expert dataset. In contrast, Ghasemipour et al. (2019) showed that adversarial imitation learning (AIL) methods, such as GAIL (Ho and Ermon, 2016), leverage a recovered reward function to perform policy optimization on “out-of-expert-data” through online interaction, significantly improving performance. Following the formulation in (Xu et al., 2022), the training objective of reward-model-based AIL can be re-formulated as
$\pi_{\operatorname{AIL}}\leftarrow\mathop{\rm argmin}_{\pi}\sum_{s\in\mathcal{S}}\sum_{a\in\mathcal{A}}\left|d_{\pi}(s,a)-\widehat{d}_{\operatorname{E}}(s,a)\right|,$
where $\widehat{d}_{\operatorname{E}}$ is the empirical state-action distribution estimated from $D_{\operatorname{E}}$, and $d_{\pi}(s,a)$ is obtained from online interaction. For the above optimization problem, states beyond those in the expert dataset are used. We notice that Xu et al. (2022) theoretically proved that AIL, by leveraging additional online data for imitation, can outperform BC.
Additionally, our research is related to transition-model-based reinforcement learning (RL) methods, where the goal is to find an optimal policy through interactions with environments. Many empirical successes suggest that transition-model-based approaches are superior in terms of sample complexity (Luo et al., 2019; Janner et al., 2019). We do not aim to present a detailed discussion since RL involves lots of concepts and notations. Instead, we would like to highlight that our findings align with the understanding that additional policy optimization on transition-model-generated data is helpful. We would like to refer readers to (Hafner et al., 2020; Schrittwieser et al., 2020; Yu et al., 2020; Luo et al., 2023) for the effect of data augmentation in transition-model-based RL methods.
Finally, we note that compared with reward-model-free methods such as DPO (Rafailov et al., 2023), reward-model-based policy optimization (RMB-PO) methods do not require extra preference annotation. For applications such as large language models (LLMs), training and storing a reward model has been shown to be highly efficient, as demonstrated in (Yao et al., 2023). The primary challenge in RMB-PO lies in the huge action space during policy optimization. However, this issue can be effectively addressed by computationally efficient methods like those proposed by (Dong et al., 2023; Li et al., 2023).
5 Conclusion
Our results underscore the importance of policy optimization on out-of-preference data to unlock the reward’s generalization capacity. Otherwise, the policy may conflict with the learned reward model on out-of-preference data and suffer a poor performance.

### 25 Feb 2024 v2

@@
 Abstract
-Aligning intelligent agents with human preferences and values is important.
-This paper examines two popular alignment methods: Direct Preference Optimization (DPO) and Reward-Model-Based Policy Optimization (RMB-PO).
-A variant of RMB-PO, referred to as RMB-PO+ is also considered.
-These methods, either explicitly or implicitly, learn a reward model from preference data and differ in the data used for policy optimization to unlock the generalization ability of the reward model.
-In particular, compared with DPO, RMB-PO additionally uses policy-generated data, and RMB-PO+ further leverages new, preference-free data.
-We examine the impact of such out-of-preference data.
-Our study, conducted through controlled and synthetic experiments, demonstrates that DPO performs poorly, whereas RMB-PO+ performs the best.
-In particular, even when providing the policy model with a good feature representation, we find that policy optimization with adequate out-of-preference data significantly improves performance by harnessing the reward model’s generalization capabilities.
-[^1] Code is available at https://github.com/liziniu/policy_optimization
+Aligning agents with human preferences is important.
+This paper examines two classes of alignment methods.
+The first class operates without explicitly learning a reward model from preference data, with Direct Preference Optimization (Rafailov et al., 2023) emerging as a prominent method within this class.
+The second class involves methods that explicitly learn a reward function and utilize it to optimize policy on prompts-only data, with Proximal Policy Optimization (Schulman et al., 2017) standing out as a popular choice.
+Within this class, we investigate a notable approach that leverages a large amount of prompts, extending beyond those present in the preference dataset.
+Experiments demonstrate that this approach outperforms other methods on synthetic contextual bandits, which serve as mathematical models for alignment.
+Additionally, we provide an analysis of source errors in these optimization methods and draw connections with other related research areas, such as imitation learning and reinforcement learning.
+In essence, our research highlights the importance of integrating out-of-preference data, including the policy’s responses to prompts from the preference dataset and new prompts, into the policy optimization.
+[^1] A short version of this paper is presented at the tiny paper track of the 12th International Conference on Learning Representations (ICLR), 2024.
+Code is available at https://github.com/liziniu/policy_optimization.
 1 Introduction
-Developing intelligent agents requires alignment with human values (Russell and Norvig, 2010).
+Developing trustworthy agents requires alignment with human preferences (Russell and Norvig, 2010).
 A standard practice involves providing a human preference dataset for the agent to learn from.
 According to utility theory (Fishburn et al., 1979), preference is connected with a certain reward function.
-Currently, there are two kinds of alignment methods: the reward-model-free approach, such as Direct Preference Optimization (DPO) (Rafailov et al., 2023) and Identity Policy Optimization (IPO) (Azar et al., 2023), and the reward-model-based approaches, exemplified by the so-called Reinforcement Learning from Human Feedback (RLHF) (Christiano et al., 2017; Stiennon et al., 2020; OpenAI, 2023).
-Specifically, the first method directly learns a decision policy from a preference dataset, without the need to train a separate reward model.
-On the other hand, the second approach involves training a reward model from the preference data and subsequently employing policy optimization (e.g., reinforcement learning) on new, preference-free data.
-While both approaches have been shown to improve the decision policy by leveraging preference data, the superiority of one method over the other remains unclear, a determination crucial for future advancements.
-We briefly explain why this question is hard to answer.
-The main difficulty lies in the intuition that the policy model, powered by a (pre-trained) neural network or learned from preference-data, possesses a certain generalization ability.
-Therefore, it is unclear whether policy optimization on additional data, with an imperfect yet generalizable reward model, is necessary.
-In this paper, we explore the above question through controlled synthetic experiments.
-We study two contextual bandit tasks with linear function approximation and neural function approximation, respectively.
-Our studies take a stochastic optimization viewpoint.
-Concretely, we argue that the policy optimization in these methods corresponds to different versions of stochastic approximation of the (expected) reward maximization problem.
-In particular, with a reward model, stochastic approximation can be performed more accurately with out-of-preference data, which interestingly does not increase the sample complexity of preference data.
-Our experiments validate the above ideas.
-One main experiment, where we manually ensure that the policy shares the same good feature representation with the reward model, shows that policy optimization with additional out-of-preference data still improves generalization performance.
+Currently, there are two kinds of alignment methods:
+  • The first class of methods, referred to as the reward-model-free approach in this paper, does not explicitly learn a reward model but directly optimizes the language model from preference annotations.
+This class includes popular algorithms such as Direct Preference Optimization (DPO) (Rafailov et al., 2023) and Identity Policy Optimization (IPO) (Azar et al., 2023).
+  • The second class of methods, referred to as the reward-model-based approaches in this paper, is exemplified by the so-called Reinforcement Learning from Human Feedback (RLHF) framework (Christiano et al., 2017; Stiennon et al., 2020; OpenAI, 2023) framework, with Proximal Policy Optimization (PPO) (Schulman et al., 2017) standing out as a popular choice.
+In particular, this class of methods trains a reward model from the preference data and subsequently optimizes the language model to improve its responses to prompts.
+While both approaches are able to improve performance by leveraging preference data, the superiority of one method over the other remains an open question, crucial for driving future advancements.
+We briefly explain the challenges in determining this superiority.
+First, we expect that the reward model, learned from preference data, possesses a certain generalization capability (i.e., through fine-tuning a powerful pre-trained neural network).
+To improve the language model, we require prompt-response pairs (i.e., input-output pairs) and their associated reward values (i.e., the supervision signals).
+This raises an important question: how should we select these prompt-response pairs?
+Notably, reward-model-free approaches utilize previously collected prompt-response pairs from the preference dataset, whereas reward-model-based approaches generate new responses to prompts using the language model, discarding the responses in the preference dataset.
+Which approach is more effective?
+And how do these data sources impact the generalization performance?
+We explore the above questions by analyzing the errors in the language model’s optimization under the framework of contextual bandits (Lattimore and Szepesvári, 2020), which serve as mathematical models for alignment.
+This analysis helps us predict the algorithm’s behaviors without the need for extensive experiments.
+Within this context, the language model is framed as a “policy” in broader terms.
+We demonstrate that the policy optimization in these methods corresponds to various forms of Monte Carlo approximations for maximizing the expected reward.
+Notably, the inclusion of out-of-preference data, which includes responses to prompts from the preference dataset and responses to new prompts, enhances the accuracy of the Monte Carlo approximation.
+To validate the above ideas, we conduct experiments on contextual bandits with linear function approximation and neural function approximation, respectively.
+One main experiment, where we manually ensure that the policy shares the same good feature representation with the reward model (thus, they have the same representation power), shows that policy optimization with additional out-of-preference data still improves generalization performance.
 Other experiments also support this claim.
-Finally, we provide a discussion about this phenomenon with reference to other fields, such as imitation learning and reinforcement learning.
+Finally, we provide a discussion about this phenomenon with reference to other fields, such as imitation learning (Osa et al., 2018) and reinforcement learning (Sutton and Barto, 2018).
 2 Problem Formulation
-We consider the so-called contextual bandit (Langford and Zhang, 2007; Lu et al., 2010) formulation for the alignment problem.
-Let $s$ and $a$ be the state (i.e., context) and decision action, respectively.
+We consider the so-called contextual bandits (Langford and Zhang, 2007; Lu et al., 2010) formulation, which serves mathematical models for alignment.
+Let $s$ and $a$ be the state and action, respectively.
 We aim to obtain a decision policy $\pi$ that acts optimally in terms of reward maximization:
-$\pi^{\star}\leftarrow\mathop{\rm argmax}_{\pi}\mathbb{E}_{s\sim\rho(\cdot)}\mathbb{E}_{a\sim\pi(\cdot|s)}\left[r(s,a)\right],$ (1)
+$\pi^{\star}_{r}\leftarrow\mathop{\rm argmax}_{\pi}\mathbb{E}_{s\sim\rho(\cdot)}\mathbb{E}_{a\sim\pi(\cdot|s)}\left[r(s,a)\right],$ (1)
 where the symbol $\rho$ denotes the state distribution, and $r$ is the ground truth reward function.
+We omit the subscript $r$ in $\pi^{\star}_{r}$ when the context is clear.
+For language models, the term “states” refers to prompts, while “actions” denote responses.
+The language model functions as the decision-making policy.
+It is worth noting that terminologies may be used interchangeably.
 In the context of alignment, the difficulty is that the reward function is unknown but only preferences over two actions are observed.
-Typically, the Bradley-Terry-Luce (Bradley and Terry, 1952) assumption is used:
+Typically, the Bradley-Terry assumption (Bradley and Terry, 1952) is used:
 ${\mathbb{P}}\left(a>a^{\prime}|s\right)=\frac{\exp(r(s,a))}{\exp(r(s,a))+\exp(r(s,a^{\prime}))},$
 where the symbol $a>a^{\prime}$ means that $a$ is more preferred compared with $a^{\prime}$.
-Given a preference dataset $D_{\operatorname{pref}}=\{(s_{i},a_{i},a_{i}^{\prime})\}_{i=1}^{n}$, where $a_{i}>a_{i}^{\prime}$ is assumed without loss of generality, the reward learning objective is
+Given a preference dataset $D_{\operatorname{pref}}=\{(s_{i},a_{i},a_{i}^{\prime})\}_{i=1}^{n}$, where $a_{i}>a_{i}^{\prime}$ is assumed without loss of generality, the reward learning objective, derived via maximum likelihood estimation, is
 $\widehat{r}\leftarrow\mathop{\rm argmax}_{r}\sum_{i=1}^{n}\log\left(\sigma\left(r(s_{i},a_{i})-r(s_{i},a_{i}^{\prime})\right)\right),\quad(s_{i},a_{i},a_{i}^{\prime})\sim D_{\operatorname{pref}}.$ (2)
 where $\sigma(\cdot)$ is the sigmoid function.
-Let $\pi_{\operatorname{ref}}$ be a reference policy model and $\beta>0$ be a hyper-parameter.
-The Reward-Model-Based Policy Optimization (RMB-PO) approach optimizes the policy on the policy-generated data by sampling states from the preference dataset:
-$\widehat{\pi}_{\operatorname{RMB-PO}}\leftarrow\mathop{\rm argmax}_{\pi}\sum_{i=1}^{n}\mathbb{E}_{a\sim\pi(\cdot|s_{i})}\left[\widehat{r}(s_{i},a)\right]+\beta D_{\mathrm{KL}}(\pi(\cdot|s_{i}),\pi_{\operatorname{ref}}(\cdot|s_{i})),\quad s_{i}\sim D_{\operatorname{pref}}.$ (3)
-In (Ouyang et al., 2022; Touvron et al., 2023), a variant of RMB-PO, refereed to as RMB-PO+ in this paper, further leverages a new, preference-free dataset $D_{\operatorname{new}}=\{s_{j}\}_{j=1}^{m}$ for policy optimization:
-$\widehat{\pi}_{\operatorname{RMB-PO+}}\leftarrow\mathop{\rm argmax}_{\pi}\sum_{j=1}^{m}\mathbb{E}_{a\sim\pi(\cdot|s_{j})}\left[\widehat{r}(s_{j},a)\right]+\beta D_{\mathrm{KL}}(\pi(\cdot|s_{j}),\pi_{\operatorname{ref}}(\cdot|s_{j})),\,\,s_{j}\sim D_{\operatorname{new}}.$ (4)
-Note that the dataset $D_{\operatorname{new}}$ is cheap to obtain and usually $m\geq n$ (Ouyang et al., 2022).
-In addition to the above methods, the Direct Preference Optimization (DPO) is developed in (Rafailov et al., 2023), which combines Equation 2 and Equation 3 to a single supervised objective:
+This objective encourages the reward function to give a high score for the positively preferred data $(s,a)$ and a low score for the negative preferred data $(s,a^{\prime})$.
+Let $\pi_{\operatorname{ref}}$ be a reference policy model and $\beta>0$ be a hyper-parameter. Ideally, we may want to optimize the policy with this recovered reward function in population:
+$\pi^{\star}_{\widehat{r}}\leftarrow\mathop{\rm argmax}_{\pi}\mathbb{E}_{s\sim\rho(\cdot)}\left\{\mathbb{E}_{a\sim\pi(\cdot|s)}\left[\widehat{r}(s,a)\right]-\beta D_{\mathrm{KL}}(\pi(\cdot|s),\pi_{\operatorname{ref}}(\cdot|s))\right\}.$
+Here, the Kullback–Leibler (KL) penalty aims to mitigate the reward hacking and over-optimization issue (Gao et al., 2023).
+We remark that, in practice, we do not know the distribution $\rho(\cdot)$ and typically employ Monte Carlo approximations.
+That is, we use finite samples to approximate the population distribution and its expectation.
+There are two kinds of practical approaches, which we elaborate on below.
+Please also see Figure 1.
+(a) Illustration for reward-model-free approaches.
+(b) Illustration for reward-model-based approaches.
+Figure 1: Illustration for policy optimization methods. For reward-model-based approaches, the reward model learning procedure is not plotted for ease of presentation.
+Reward-model-free Optimization Approaches: One direct idea is to use state-action pairs from the preference dataset for policy optimization:
+$\widehat{\pi}_{\operatorname{RMF-PO}}\leftarrow\mathop{\rm argmax}_{\pi}\sum_{i=1}^{n}\sum_{a\in\{a_{i},a_{i}^{\prime}\}}\widehat{r}(s_{i},a)-\beta D_{\mathrm{KL}}\left(\pi(a|s_{i}),\pi_{\operatorname{ref}}(a|s_{i})\right),\quad(s_{i},a_{i},a_{i}^{\prime})\sim D_{\operatorname{pref}}.$ (3)
+This approach approximate the state and action distributions by finite samples from the preference dataset.
+By using tools from KL-regularized optimization (see e.g., (Vieillard et al., 2020)), Rafailov et al.
+(2023) showed that procedures in Equation 2 and Equation 3 could be integrated into a single objective:
 $\widehat{\pi}_{\operatorname{DPO}}\leftarrow\mathop{\rm argmax}_{\pi}\sum_{i=1}^{n}\log\sigma\left(\beta\log\frac{\pi(a_{i}|s_{i})}{\pi_{\operatorname{ref}}(a_{i}|s_{i})}-\beta\log\frac{\pi(a_{i}^{\prime}|s_{i})}{\pi_{\operatorname{ref}}(a_{i}^{\prime}|s_{i})}\right),\quad(s_{i},a_{i},a_{i}^{\prime})\sim D_{\operatorname{pref}}.$ (4)
-Note that DPO uses solely the preference data, without policy-generated data or preference-free data.
-2.1 A Stochastic Optimization Perspective
-Based on the above formulations, we argue that all the mentioned methods are stochastic approximations to Equation 1, with errors arising from three sources:
+The resultant algorithm is named Direct Preference Optimization (DPO). Since this approach does not require explicitly training a reward model, it is considered as reward-model-free optimization.
+Reward-model-based Optimization Approaches: The vanilla Reward-Model-Based Policy Optimization (RMB-PO) approach also leverages states from the preference dataset but samples actions from the policy model:
+$\widehat{\pi}_{\operatorname{RMB-PO}}\leftarrow\mathop{\rm argmax}_{\pi}\sum_{i=1}^{n}\mathbb{E}_{a\sim\pi(\cdot|s_{i})}\left[\widehat{r}(s_{i},a)\right]-\beta D_{\mathrm{KL}}(\pi(\cdot|s_{i}),\pi_{\operatorname{ref}}(\cdot|s_{i})),\quad s_{i}\sim D_{\operatorname{pref}}.$ (5)
+We consider the exact action expectation $\mathbb{E}_{a\sim\pi(\cdot|s)}[\widehat{r}(s,a)]$ in the above formulation, and this expectation can be approximated by sampling multiple actions.
+This approximation error can be mitigated by computational power, and we do not consider this error in this paper.
+In (Ouyang et al., 2022; Touvron et al., 2023), a variant of RMB-PO, refereed to as RMB-PO+ in this paper, further leverages a new, preference-free dataset $D_{\operatorname{pref-free}}=\{s_{j}\}_{j=1}^{m}$:
+$\widehat{\pi}_{\operatorname{RMB-PO+}}\leftarrow\mathop{\rm argmax}_{\pi}\sum_{j=1}^{m}\mathbb{E}_{a\sim\pi(\cdot|s_{j})}\left[\widehat{r}(s_{j},a)\right]-\beta D_{\mathrm{KL}}(\pi(\cdot|s_{j}),\pi_{\operatorname{ref}}(\cdot|s_{j})),\,\,s_{j}\sim D_{\operatorname{pref-free}}.$ (6)
+Note that the dataset $D_{\operatorname{pref-free}}$ is cheap to obtain and usually $m\geq n$ (Ouyang et al., 2022).
+One particular example of such data in language model’s application is the lmsys-chat-1m dataset (Zheng et al., 2023), which has 1 million prompts from real users without preference annotations.
+We note that there is no single learning objective for reward-model-based approaches.
+This is because the technique in (Rafailov et al., 2023) requires that the reward and policy learning objectives have the same training distribution, a condition that is not met for reward-model-based approaches.
+In practice, policy optimization in Equation 5 and Equation 6 can be conducted by policy gradient methods such as PPO (Schulman et al., 2017) and ReMax (Li et al., 2023).
+2.1 Theoretical Analysis
+In this section, we present a preliminary analysis of errors in the optimization methods. At a high level, we identify three types of errors:
   • 1)
-    the reward model error $|\widehat{r}(s,a)-r(s,a)|$;
+    the reward evaluation error $|\widehat{r}(s,a)-r(s,a)|$;
   • 2)
@@
     the estimation error when using finite samples to calculate the expectation $\mathbb{E}_{s\sim\rho(\cdot)}[\cdot]$.
-The first error primarily results from having only finite preference data, which is expensive to obtain.
+The first error primarily results from finite preference data and diminishes to zero as the preference data size increases indefinitely.
+This error exists in all optimization methods.
 Compared with DPO, RMB-PO aims to mitigate the second error, while RMB-PO+ further reduces the third error.
-We note that RMB-PO and RMB-PO+ do not increase the sample complexity of preference data but only incur additional computation steps.
-In our study, we assume that the action space $\mathcal{A}$ is finite and the expectation $\mathbb{E}_{a\sim\pi(\cdot|s)}[\cdot]$ is easy to obtain.
-For applications with combinatorial action spaces (e.g., large language models), computing this expectation is generally intractable, and practical algorithms may employ the policy to generate stochastic trials to approximate the expectation; refer to, e.g., (Li et al., 2023).
-We note that even in this case, the sample complexity of preference data does not increase.
-Another viewpoint is that the RMB-PO and RMB-PO+ methods incorporate additional policy-generated data and preference-free data, essentially employing a form of data augmentation.
-We summarize the effect of such data augmentation in Table 1.
-DPO RMB-PO RMB-PO+
-Improvement of Stochastic Approximation in State Space
-Improvement of Stochastic Approximation in Action Space
-Table 1: Illustration of the data augmentation effect in RMB-PO and RMB-PO+.
+We note that RMB-PO and RMB-PO+ do not increase the sample complexity of preference data but only incur additional computational steps.
+We present the error bound analysis below.
+Proposition 1.
+Define the reward evaluation error $\varepsilon_{r}=\sup_{(s,a)}|\widehat{r}(s,a)-r(s,a)|$, the state distribution estimation error $\varepsilon_{s}=\sup_{\pi}\sup_{r:{\mathcal{S}}\times{\mathcal{A}}\rightarrow[0,1]}\mathbb{E}_{s}\mathbb{E}_{a\sim\pi(\cdot|s)}[r(s,a)]-\widehat{\mathbb{E}}_{s}\mathbb{E}_{a\sim\pi(\cdot|s)}[r(s,a)]$, and the action distribution estimation error $\varepsilon_{a}=\sup_{\pi}\sup_{r:{\mathcal{S}}\times{\mathcal{A}}\rightarrow[0,1]}\sup_{s}\mathbb{E}_{a\sim\pi(\cdot|s)}[r(s,a)]-\widehat{\mathbb{E}}_{a\sim\pi(\cdot|s)}[r(s,a)]$.
+Here $\widehat{\mathbb{E}}_{s}$ and $\widehat{\mathbb{E}}_{a}$ denote the finite-sample estimations of expectation under the state and action distributions, respectively.
+Consider $\beta=0$, then we have
+$r(\pi^{\star}_{r})-r(\widehat{\pi}_{\operatorname{RMB-PO}})$ $\leq 2\varepsilon_{r}+2\varepsilon_{s},$
+$r(\pi^{\star}_{r})-r(\widehat{\pi}_{\operatorname{RMF-PO}})$ $\leq 2\varepsilon_{r}+2\varepsilon_{s}+2\varepsilon_{a},$
+where $r(\pi)=\mathbb{E}_{s}\mathbb{E}_{a\sim\pi(\cdot|s)}[r(s,a)]$ is the evaluation performance of a policy $\pi$.
+Proof.
+We first consider the reward error:
+$\quad r(\pi^{\star}_{r})-r(\widehat{\pi}_{\operatorname{RMB-PO}})$
+$=\mathbb{E}_{s}\mathbb{E}_{a\sim\pi^{\star}_{r}}[r(s,a)]-\mathbb{E}_{s}\mathbb{E}_{a\sim\widehat{\pi}_{\operatorname{RMB-PO}}}[r(s,a)]$
+$=\mathbb{E}_{s}\mathbb{E}_{a\sim\pi^{\star}_{r}}[r(s,a)]-\mathbb{E}_{s}\mathbb{E}_{a\sim\pi^{\star}_{r}}[\widehat{r}(s,a)]+\mathbb{E}_{s}\mathbb{E}_{a\sim\pi^{\star}_{r}}[\widehat{r}(s,a)]-\mathbb{E}_{s}\mathbb{E}_{a\sim\widehat{\pi}_{\operatorname{RMB-PO}}}[r(s,a)]$
+$\leq\varepsilon_{r}+\mathbb{E}_{s}\mathbb{E}_{a\sim\pi^{\star}_{r}}[\widehat{r}(s,a)]-\mathbb{E}_{s}\mathbb{E}_{a\sim\widehat{\pi}_{\operatorname{RMB-PO}}}[\widehat{r}(s,a)]+\mathbb{E}_{s}\mathbb{E}_{a\sim\widehat{\pi}_{\operatorname{RMB-PO}}}[\widehat{r}(s,a)]-\mathbb{E}_{s}\mathbb{E}_{a\sim\widehat{\pi}_{\operatorname{RMB-PO}}}[r(s,a)]$
+$\leq 2\varepsilon_{r}+\mathbb{E}_{s}\mathbb{E}_{a\sim\pi^{\star}_{r}}[\widehat{r}(s,a)]-\mathbb{E}_{s}\mathbb{E}_{a\sim\widehat{\pi}_{\operatorname{RMB-PO}}}[\widehat{r}(s,a)].$ (7)
+Then we consider the state distribution estimation error:
+$\quad\mathbb{E}_{s}\mathbb{E}_{a\sim\pi^{\star}_{r}}[\widehat{r}(s,a)]-\mathbb{E}_{s}\mathbb{E}_{a\sim\widehat{\pi}_{\operatorname{RMB-PO}}}[\widehat{r}(s,a)]$
+$=\mathbb{E}_{s}\mathbb{E}_{a\sim\pi^{\star}_{r}}[\widehat{r}(s,a)]-\widehat{\mathbb{E}}_{s}\mathbb{E}_{a\sim\pi^{\star}_{r}}[\widehat{r}(s,a)]+\widehat{\mathbb{E}}_{s}\mathbb{E}_{a\sim\pi^{\star}_{r}}[\widehat{r}(s,a)]-\mathbb{E}_{s}\mathbb{E}_{a\sim\widehat{\pi}_{\operatorname{RMB-PO}}}[\widehat{r}(s,a)]$
+$\leq\varepsilon_{s}+\widehat{\mathbb{E}}_{s}\mathbb{E}_{a\sim\pi^{\star}_{r}}[\widehat{r}(s,a)]-\widehat{\mathbb{E}}_{s}\mathbb{E}_{a\sim\widehat{\pi}_{\operatorname{RMB-PO}}}[\widehat{r}(s,a)]+\widehat{\mathbb{E}}_{s}\mathbb{E}_{a\sim\widehat{\pi}_{\operatorname{RMB-PO}}}[\widehat{r}(s,a)]-\mathbb{E}_{s}\mathbb{E}_{a\sim\widehat{\pi}_{\operatorname{RMB-PO}}}[\widehat{r}(s,a)]$
+$\leq\varepsilon_{s}+0+\varepsilon_{s}.$ (8)
+Combining (7) and (8) proves the first result in Proposition 1.
+For the second result, we may replace $\widehat{\pi}_{\operatorname{RMB-PO}}$ with $\widehat{\pi}_{\operatorname{RMF-PO}}$ in the above proof to obtain:
+$r(\pi^{\star}_{r})-r(\widehat{\pi}_{\operatorname{RMF-PO}})\leq 2\varepsilon_{r}+2\varepsilon_{s}+\widehat{\mathbb{E}}_{s}\mathbb{E}_{a\sim\pi^{\star}_{r}}[\widehat{r}(s,a)]-\widehat{\mathbb{E}}_{s}\mathbb{E}_{a\sim\widehat{\pi}_{\operatorname{RMF-PO}}}[\widehat{r}(s,a)].$ (9)
+Then we consider the action distribution estimation error:
+$\quad\widehat{\mathbb{E}}_{s}\mathbb{E}_{a\sim\pi^{\star}_{r}}[\widehat{r}(s,a)]-\widehat{\mathbb{E}}_{s}\mathbb{E}_{a\sim\widehat{\pi}_{\operatorname{RMF-PO}}}[\widehat{r}(s,a)]$
+$=\widehat{\mathbb{E}}_{s}\mathbb{E}_{a\sim\pi^{\star}_{r}}[\widehat{r}(s,a)]-\widehat{\mathbb{E}}_{s}\widehat{\mathbb{E}}_{a\sim{\pi}^{\star}_{r}}[\widehat{r}(s,a)]+\widehat{\mathbb{E}}_{s}\widehat{\mathbb{E}}_{a\sim{\pi}^{\star}_{r}}[\widehat{r}(s,a)]-\widehat{\mathbb{E}}_{s}\mathbb{E}_{a\sim\widehat{\pi}_{\operatorname{RMF-PO}}}[\widehat{r}(s,a)]$
+$\leq\varepsilon_{a}+\widehat{\mathbb{E}}_{s}\widehat{\mathbb{E}}_{a\sim{\pi}^{\star}_{r}}[\widehat{r}(s,a)]-\widehat{\mathbb{E}}_{s}\widehat{\mathbb{E}}_{a\sim\widehat{\pi}_{\operatorname{RMF-PO}}}[\widehat{r}(s,a)]+\widehat{\mathbb{E}}_{s}\widehat{\mathbb{E}}_{a\sim\widehat{\pi}_{\operatorname{RMF-PO}}}[\widehat{r}(s,a)]-\widehat{\mathbb{E}}_{s}\mathbb{E}_{a\sim\widehat{\pi}_{\operatorname{RMF-PO}}}[\widehat{r}(s,a)]$
+$\leq\varepsilon_{a}+0+\varepsilon_{a}$ (10)
+Combining (9) and (10) proves the second result in Proposition 1. ∎
+We do not present the analysis for $\widehat{\pi}_{\operatorname{RMB-PO+}}$, as its analysis is similar to that for $\widehat{\pi}_{\operatorname{RMB-PO}}$.
+The main difference lies in that the state distribution estimation error for $\widehat{\pi}_{\operatorname{RMB-PO+}}$ is generally smaller than that for $\widehat{\pi}_{\operatorname{RMB-PO}}$ due to more samples.
+We note that our analysis is quite basic in the sense that we consider all errors in the supremum norm, and it would be interesting to explore a more tighter analysis with finite sample guarantee; see e.g., (Xiong et al., 2023) for recent progress in this direction.
 3 Experiments
 In this section, we conduct numerical experiments to validate the improvement of RMB-PO and RMB-PO+ by better stochastic approximation.
-All of our experiments are run with 10 different random seeds (2021-2030), and the averaged results are reported.
+All of our experiments are run with 10 different random seeds (2021-2030), and the averaged results are reported [^2] We exclude the worst and best results to make a robust estimation of the performance..
 Note that we set $\pi_{\text{REF}}$ to be a policy with a uniform action distribution in all experiments and $\beta=0.01$ for all methods.
@@
 $\phi_{\pi}(s,a)=\left((a+1)\cdot\sin(s\cdot\pi),\frac{1}{a+1}\cdot\cos(s\cdot\pi)\right)^{\top}.$
-In our experiments, we set the size of preference data to be $n=20$ and the size of preference-free data to be $m=10n$, resulting in training accuracy of the reward model ranging from 60% to 80%.
-We display the optimality gap $|r(\pi^{\star})-r(\widehat{\pi})|$ (the smaller, the better) in Figure 2 and Figure 2, where $r(\pi)$ is the expected performance, i.e., $r(\pi)=\mathbb{E}_{s\sim\rho(\cdot)}\mathbb{E}_{a\sim\pi(\cdot|s)}[r(s,a)]$ (in our experiments, we use 5000 sampled states to approximate this expectation).
+We believe that in scenarios where $\phi_{\pi}\neq\phi_{r}$, RMB-PO approaches could exhibit more promising performance than RMF-PO approaches.
+This is because, in such cases, the policy and reward models may align well by learning from preference data.
+However, in out-of-preference-distribution scenarios, they may extrapolate and generalize quite differently due to mismatches in representations.
+Nevertheless, RMB-PO approaches could use out-of-preference-distribution data to mitigate these mismatches and tend to perform well.
+The case where $\phi_{\pi}\neq\phi_{r}$ will be revisited in later neural bandit experiments, where the policy model and reward model typically utilize distinct architectures and learn distinct representations.
+In our experiments, we set the size of preference data to be $n=20$ and the size of preference-free data to be $m=10n$, resulting in training accuracy of the reward model ranging from 60% to 80% over 10 experiments.
+We display the optimality gap $|r(\pi^{\star})-r(\widehat{\pi})|$ (the smaller, the better) in Figure 3 and Figure 3, where $r(\pi)$ is the evaluation performance of a policy $\pi$, i.e., $r(\pi)=\mathbb{E}_{s\sim\rho(\cdot)}\mathbb{E}_{a\sim\pi(\cdot|s)}[r(s,a)]$ (in our experiments, we use 5000 sampled states to approximate this expectation).
 From Figure 3, we see that even though the policy model is provided with a good feature (e.g., in Figure 3), RMB-PO methods can benefit from out-of-preference data.
@@
 Thus, we believe it is crucial to learn the optimal action (as inferred by the reward model) on out-of-preference data, even when the two models share the same good feature.
+To gain a better understanding, we also visualize the learned policy distribution in the $\phi_{\pi}\neq\phi_{r}$ setting; see Figure 4.
+To observe the training distribution coverage, we plot the states from the preference dataset.
+Additional states used in RMB-PO+ almost cover the entire state space but are not shown for readability reasons.
+From the reported curves, we observe that DPO aligns well with the optimal policy in the regions covered by preference data, and RMB-PO(+) methods tend to perform better than DPO in the out-of-distribution regime not covered by the preference data.
+(a) Action $a_{0}$.
+(b) Action $a_{1}$.
+(c) Action $a_{2}$.
+(d) Action $a_{3}$.
+Figure 4: Probabilities of four actions $a_{0}$, $a_{1}$, $a_{2}$ and $a_{3}$.
+Results illustrate that RMB-PO(+) methods leverage out-of-preference data to better learn the policy distribution on out-of-distribution states and improve the generalization performance.
 Following the same setup, we provide ablation studies regarding the size of preference-free data used in RMB-PO+.
@@
 4 Discussion
-Our results are related to imitation learning (Osa et al., 2018), where the goal is to learn a policy from expert demonstrations.
+Our research is related to imitation learning (Osa et al., 2018), which aims to learn a policy from expert demonstrations.
 A popular approach to achieve this goal is through behavioral cloning (BC) (Pomerleau, 1991), which trains a policy model by maximizing the likelihood of expert data.
@@
 where $D_{\operatorname{E}}$ is the expert dataset.
-In contrast, Ghasemipour et al.
-(2019) showed that adversarial imitation learning (AIL) methods, such as GAIL (Ho and Ermon, 2016), leverage a recovered reward function to perform policy optimization on “out-of-expert-data” through online interaction, significantly improving performance.
+Ghasemipour et al.
+(2019) showed that another class of imitation methods, known as adversarial imitation learning (AIL) methods, (such as GAIL (Ho and Ermon, 2016)), usually performs better than BC.
+In particular, AIL methods leverage a recovered reward function to perform policy optimization on “out-of-expert-data” through online interaction, significantly improving performance.
 Following the formulation in (Xu et al., 2022), the training objective of reward-model-based AIL can be re-formulated as
@@
 where $\widehat{d}_{\operatorname{E}}$ is the empirical state-action distribution estimated from $D_{\operatorname{E}}$, and $d_{\pi}(s,a)$ is obtained from online interaction.
-For the above optimization problem, states beyond those in the expert dataset are used.
+For the optimization objective of AIL, it utilizes states beyond those in the expert dataset (reflected in the summation over all state-action pairs).
 We notice that Xu et al.
-(2022) theoretically proved that AIL, by leveraging additional online data for imitation, can outperform BC.
+(2022) theoretically proved that AIL can outperform BC in terms of addressing the distribution shift issue with optimization on “out-of-expert data”.
+The idea of recovering a reward function and using it to perform extensive policy optimization is quite similar to the framework of RLHF.
 Additionally, our research is related to transition-model-based reinforcement learning (RL) methods, where the goal is to find an optimal policy through interactions with environments.
@@
 Finally, we note that compared with reward-model-free methods such as DPO (Rafailov et al., 2023), reward-model-based policy optimization (RMB-PO) methods do not require extra preference annotation.
-For applications such as large language models (LLMs), training and storing a reward model has been shown to be highly efficient, as demonstrated in (Yao et al., 2023).
+For applications such as language models, training and storing a reward model has been shown to be highly efficient, as demonstrated in (Yao et al., 2023).
 The primary challenge in RMB-PO lies in the huge action space during policy optimization.
 However, this issue can be effectively addressed by computationally efficient methods like those proposed by (Dong et al., 2023; Li et al., 2023).
+Notably, Li et al.
+(2023) showed that optimizing the language model with prompts-only data can improve performance, a setting that cannot achieved by reward-model-free approaches such as DPO.
 5 Conclusion
-Our results underscore the importance of policy optimization on out-of-preference data to unlock the reward’s generalization capacity.
-Otherwise, the policy may conflict with the learned reward model on out-of-preference data and suffer a poor performance.
+We analyze the errors of policy optimization methods when learning from preferences for alignment.
+We also conduct experiments to validate our claims.
+Our results underscore the importance of optimizing policies on out-of-preference data and the power of using a reward model to provide supervision signals.